# Efficient and Flexible Multirate Temporal Adaptivity

### v1

@@ Abstract
In this work we present two new families of multirate time step adaptivity controllers, that are designed to work with embedded multirate infinitesimal (MRI) time integration methods for adapting time steps when solving problems with multiple time scales. We compare these controllers against competing approaches on two benchmark problems and see that they offer dramatically improved performance and flexibility, with each proposed family excelling on different types of multirate applications. The combination of embedded MRI methods and the proposed controllers enable adaptive simulations of problems with a potentially arbitrary number of time scales, achieving high accuracy while maintaining low computational cost. Additionally, we introduce a new set of embeddings for the family of explicit multirate exponential Runge–Kutta (MERK) methods of orders 2 through 5, resulting in the first-ever fifth-order embedded MRI method. Finally, we compare the performance of a wide range of embedded MRI methods on our benchmark problems to provide guidance on how to select an appropriate MRI method and multirate controller.
@@
journal: Applied Mathematics and Computation
@@ 1 Introduction
In this work, we focus on time-step adaptivity within multirate infinitesimal (MRI) methods, that are used when solving systems of ordinary differential equation initial value problems of the form
@@
MRI methods are determined by their fast stage time intervals $\theta_{0,i},\theta_{F,i}]$, initial conditions $v_{0,i}$, forcing functions $r_{i}(\theta)$, and embedding forcing function $\tilde{r}_{s}(\theta)$. Both $r_{i}(\theta)$ and $\tilde{r}_{s}(\theta)$ are constructed using linear combinations of $\left\{f^{\mathfrak{s}}(\theta_{F,j},Y_{j})\right\}$, and serve to propagate information from the slow to the fast time scales. Implicit and implicit-explicit extensions of the above MRI algorithm exist, that involve replacing step 2a with an implicit solve for some internal stages $Y_{i}$. The embedded solution $\tilde{y}_{n+1}$ is similar to embedded solutions in standard Runge-Kutta methods, in that it approximates the solution to an alternate order of accuracy and may be computed with minimal extra effort beyond computing $y_{n+1}$; in this case it is computed by re-solving the last stage with an alternate forcing function, $\tilde{r}_{s}(\theta)$.
As seen in steps 2a and 3 above, computation of each stage in an explicit MRI method requires solving a secondary, inner, IVP. Typically, these inner IVPs are not solved exactly, and instead are approximated using another numerical method with inner time steps $h_{n,m}$ such that $\max_{m}|h_{n,m}|\ll|H_{n}|$.
In the sections that follow, we first summarize how temporal adaptivity is performed for problems with a single time scale in Section 2. In Section 3 we introduce previous work, and propose two new approaches, for dynamically adapting both the slow and fast time steps, $H_{n}$ and $h_{n,m}$, within MRI methods. Two of these multirate controller families require estimates of the accumulated fast temporal error, so we describe our algorithms for computing those estimates in Section 4. The majority of this manuscript focuses on numerical tests of these multirate controllers in Section 5. In that section, we first describe our two benchmark problems that we use throughout our numerical tests. We then outline the set of embedded MRI methods that we will test, including new embeddings for the explicit multirate exponential Runge–Kutta (MERK) methods that result in the first ever fifth-order embedded MRI method. The remainder of Section 5 performs a variety of numerical experiments to examine the proposed multirate adaptivity strategies and the embedded MRI methods. We discuss our conclusions from these tests and identify future directions to extend this work in Section 6.
@@ 2 Single Rate Control
Before introducing multirate temporal adaptivity, we briefly review the typical approaches for single-rate adaptivity, that attempt to control the local error induced within a given time step. Assuming that the initial condition at the time step $t_{n}$ is exact, i.e., $y_{n}-y(t_{n})=0$, the local error is defined to be
@@
where $\|\cdot\|$ is a norm that incorporates the requested temporal accuracy (e.g., the WRMS-norm from [7, 8]), and (b) perform as few time steps as possible to meet objective (a). Additional considerations include a smooth step size progression [9, 10, 11, 12], and improving the efficiency and robustness of iterative algebraic solvers that may be used at each step [13, 14].
@@
    compute a candidate approximation $y_{n+1}^{*}\approx y(t_{n}+h_{n})$ and a corresponding approximation of the local error $\varepsilon_{n}^{*}\approx y(t_{n}+h_{n})-y_{n+1}^{*}$.
@@
Since the true local error $\varepsilon_{n}$ is inaccessible, then for simplicity in the remainder of this paper, we use $\varepsilon_{n}$ to denote the approximation $\varepsilon_{n}^{*}$.
Whether $\varepsilon_{n}$ passes or fails the error test (3), step (iii) must select $\tilde{h}$ to use on the next step attempt. This adaptivity controller is a function that typically depends on a small set of $(h_{n-k},\varepsilon_{n-k})$ pairs, i.e.,
$\tilde{h}=\mathcal{C}(h_{n},\varepsilon_{n},h_{n-1},\varepsilon_{n-1},\ldots,p).$
There are a myriad of choices for controllers $\mathcal{C}$ in the literature, but the simplest approach is the so-called I controller. Assuming that the numerical method for computing $y_{n+1}$ has order $p$, then from [15]
@@
for some $c(t)$ independent of $h_{n}$. Taking the norm of this gives
@@
Since we have just computed the estimate $\|\varepsilon_{n}\|$ based off a step with size $h_{n}$, a piecewise constant approximation $c(t)\approx c_{n}$ allows us to “solve” for $\|c_{n}\|=\|\varepsilon_{n}\|/h_{n}^{p+1}$. Assuming that $c(t)$ will remain relatively constant from one step attempt to the next, the candidate step size $\tilde{h}$ is computed to exactly attain an error goal of 1 (or any toltol):
$tol=\|c_{n}\|\tilde{h}^{p+1}\quad\Rightarrow\quad\tilde{h}=h_{n}\left(\frac{tol}{\|\varepsilon_{n}\|}\right)^{1/(p+1)}.$
This formula automatically grows or shrinks the step size based on whether the error test $\|\varepsilon_{n}\|^{1/(p+1)}\leq tol$ passed or failed, respectively. Due to the multiple assumptions above, a safety factor $\sigma<1$ is typically applied,
$\tilde{h}=\sigma h_{n}\left(\frac{tol}{\|\varepsilon_{n}\|}\right)^{1/(p+1)}.$
More advanced approaches for $\mathcal{C}$ are typically based on control theory, and use additional $(h_{n-k},\varepsilon_{n-k})$ values to build higher-degree piecewise polynomial approximations of the principal error function [9, 13, 14, 10, 11, 12].
@@ 3 Multirate Temporal Controllers
Multirate temporal adaptivity has not been widely studied in the literature. The two articles we are aware of are [16] and [17]. The former is specific to “MrGARK” methods [18], that require a more rigid algorithmic structure than the MRI algorithm introduced in Section 1. The latter works with MRI methods, and will serve as a baseline comparison for our new methods. In the following subsections, we consider three such approaches in this work.
@@ 3.2 Decoupled” Multirate Control
where $P$ and $p$ are the orders of the MRI method and inner solver, respectively. Here, $(H_{k},\varepsilon_{k}^{\mathfrak{s}})$ are the stepsize and local error estimates for time step $k$ at the slow time scale, and $(h_{k,\ell},\varepsilon_{k,\ell}^{\mathfrak{f}})$ are the stepsize and local error estimates for the fast substep $\ell$ within the slow step $k$. We note that in this approach, both $\mathcal{C}^{\mathfrak{s}}$ and $\mathcal{C}^{\mathfrak{f}}$ are decoupled, and thus selection of $\tilde{H}$ and $\tilde{h}$ occur independently through application of any pair of single-rate controllers. We summarize their use in the following pseudocode.
@@
    Let: $y_{n+1}=Y_{s}$, $\tilde{y}_{n+1}=\tilde{v}_{s}(\theta_{F,s})$, and $\varepsilon_{n}^{\mathfrak{s}}=y_{n+1}-\tilde{y}_{n+1}$.
@@ 3.3 Stepsize-tolerance ($H\text{-}Tol$”) Control
Our second proposed family are the so-called $H\text{-}Tol$” multirate controllers. As outlined in Section 2, standard controllers attempt to adapt step sizes to control local error within each step to achieve a requested tolerance. However, MRI methods must ask another “inner” adaptive fast-scale solver to produce the stage solutions $v_{i}(\theta_{F,i})$, that result from sub-stepping over each interval $\theta_{0,i},\theta_{F,i}]$. Local errors within the fast integrator may accumulate, resulting in an overall fast-solver error $\varepsilon^{\mathfrak{f}}_{i}$ that could exceed the requested tolerance. If that inner solver can produce both $v_{i}(\theta_{F,i})$ and an estimation of the accumulated error, $\varepsilon^{\mathfrak{f}}_{i,approx}$, then the tolerances provided to that next-fastest solver can be adjusted accordingly to ensure MRI stage solutions $Y_{i}=v_{i}(\theta_{F,i})$ that are within the overall tolerances requested of the outer MRI method.
@@
To this end, first assume that at step $n$, the MRI method requests solutions from the inner solver that are $\text{tolfac}_{n}$ more accurate than the MRI method, i.e., if the MRI method strives to achieve local errors $\|\varepsilon_{n}^{\mathfrak{s}}\|\leq 1$, then the inner solver strives to achieve local substep errors $\|\varepsilon_{n,m}^{\mathfrak{f}}\|\leq\text{tolfac}_{n}$. To account for accumulation of temporal error in the inner solver, we assume that its accumulated error over the slow step $t_{n},t_{n}+H_{n}]$ has the form
$\|\varepsilon^{\mathfrak{f}}_{n}\|=\chi(t_{n})H_{n}\text{tolfac}_{n},$ (6)
where $\chi(t)$ is independent of $\text{tolfac}_{n}$, but may vary in time. We construct this ansatz as follows: since the fast integrator takes multiple substeps to integrate over the full step $H_{n}$ then the accumulated error should be proportional to this interval width; the problem- and method-dependent factor $\chi(t_{n})$ models this constant of proportionality.
@@
By grouping the terms from (6) as $\|\varepsilon^{\mathfrak{f}}_{n}\|=\left(\chi(t_{n})H_{n}\right)\left(\text{tolfac}_{n}\right)^{0+1}$, we see that this fits the standard asymptotic error assumption used in single-rate controllers (4), through identifying $\chi(t_{n})H_{n}$ with $\|c(t_{n})\|$, the control parameter $\text{tolfac}_{n}$ with $h_{n}$, and by defining the “order” $p=0$. Thus, any single-rate controller that relies on the asymptotic error assumption (6) may be used to adjust $\text{tolfac}_{n}$ between slow step attempts.
@@
    Let: $y_{n+1}=Y_{s}$, $\tilde{y}_{n+1}=\tilde{v}_{s}(\theta_{F,s})$, $\varepsilon_{n}^{\mathfrak{s}}=y_{n+1}-\tilde{y}_{n+1}$, and retrieve the accumulated fast error $\varepsilon^{f}_{n}$ from the inner solver.

### v2

@@ Abstract
In this work we present two new families of multirate time step adaptivity controllers, that are designed to work with embedded multirate infinitesimal (MRI) time integration methods for adapting time steps when solving problems with multiple time scales. We compare these controllers against competing approaches on two benchmark problems, showing that the proposed methods offer dramatically improved performance and flexibility. The combination of embedded MRI methods and the proposed controllers enable adaptive simulations of problems with a potentially arbitrary number of time scales, achieving high accuracy while maintaining low computational cost. Additionally, we introduce a new set of embeddings for the family of explicit multirate exponential Runge–Kutta (MERK) methods of orders 2 through 5, resulting in the first-ever fifth-order embedded MRI method. Finally, we compare the performance of a wide range of embedded MRI methods on our benchmark problems to provide guidance on how to select an appropriate MRI method and multirate controller.
@@
journal: Journal of Computational and Applied Mathematics
@@ 1 Introduction
In this work, we focus on time-step adaptivity within multirate infinitesimal (MRI) methods that are used when solving systems of ordinary differential equation initial value problems (IVPs) of the form
@@
MRI methods are determined by their fast stage time intervals $\theta_{0,i},\theta_{F,i}]$, initial conditions $v_{0,i}$, forcing functions $r_{i}(\theta)$, and embedding forcing function $\tilde{r}_{s}(\theta)$. Both $r_{i}(\theta)$ and $\tilde{r}_{s}(\theta)$ are constructed using linear combinations of $\left\{f^{\mathfrak{s}}(\theta_{F,j},Y_{j})\right\}$, and serve to propagate information from the slow to the fast time scales. Implicit and implicit-explicit extensions of the above MRI algorithm exist, that involve replacing step 2a with an implicit solve for some internal stages $Y_{i}$. The embedded solution $\tilde{y}_{n+1}$ is similar to embedded solutions in standard Runge–Kutta methods, in that it approximates the solution to an alternate order of accuracy and may be computed with minimal extra effort beyond computing $y_{n+1}$; in this case it is computed by re-solving the last stage with an alternate forcing function, $\tilde{r}_{s}(\theta)$.
As seen in steps 2a and 3 above, computation of each stage in an explicit MRI method requires solving a secondary, inner, IVP. Typically, these inner IVPs are not solved exactly, and instead are approximated using another numerical method with inner time steps $h_{n,m}$ such that $\max_{m}|h_{n,m}|\ll|H_{n}|$. Our goal in this paper is to examine optimal “controller” algorithms that can adaptively select both $H_{n}$ and $h_{n,m}$ within MRI methods.
In the sections that follow, we first summarize how temporal adaptivity is performed for problems with a single time scale. In Section 3 we introduce previous work, and propose two new approaches, for dynamically adapting both the slow and fast time steps, $H_{n}$ and $h_{n,m}$, within MRI methods. Two of these multirate controller families require estimates of the accumulated fast temporal error, so we describe our algorithms for computing those estimates in Section 4. The majority of this manuscript focuses on numerical tests of these multirate controllers in Section 5. In that section, we first describe our two benchmark problems that we use throughout our numerical tests. We then outline the set of embedded MRI methods that we will test, including new embeddings for the explicit multirate exponential Runge–Kutta (MERK) methods that result in the first ever fifth-order embedded MRI method. The remainder of Section 5 performs a variety of numerical experiments to examine the proposed multirate adaptivity strategies and the embedded MRI methods. We discuss our conclusions from these tests and identify future directions to extend this work in Section 6.
@@ 2 Single Rate Control
Before introducing multirate temporal adaptivity, we briefly review the typical approaches for single-rate adaptivity, that attempt to control the local error induced within a given time step. For additional information on this topic, we refer the reader to the references [7, 8, 9]. Assuming that the initial condition at the time step $t_{n}$ is exact, i.e., $y_{n}-y(t_{n})=0$, the local error is defined to be
@@
where $\|\cdot\|$ is a norm that incorporates the requested temporal accuracy, such as the weighted root-mean-square (WRMS) norm from [8, 9],
$\|\varepsilon_{n}\|=\left(\frac{1}{N}\sum_{j=1}^{N}\left(\frac{|\varepsilon_{n,j}|}{\text{reltol}|y_{n,j}|+\text{abstol}_{j}}\right)^{2}\right)^{1/2}$ (4)
for IVPs with $N$ components, where reltol and $\text{abstol}_{j}$ are user-specified relative and absolute tolerances, respectively, and (b) perform as few time steps as possible to meet objective (a). Additional considerations include a smooth step size progression [10, 11, 12, 13], and improving the efficiency and robustness of iterative algebraic solvers that may be used at each step [14, 15].
@@
    compute a candidate approximation $y_{n+1}^{*}\approx y(t_{n}+h_{n})$ and an approximation of the local error $\varepsilon_{n}^{*}=\tilde{y}_{n+1}-y_{n+1}^{*}\approx y(t_{n}+h_{n})-y_{n+1}^{*}$.
@@
Whether $\varepsilon_{n}^{*}$ passes or fails the error test (3), step (iii) must select $\tilde{h}$ to use on the next step attempt. This adaptivity controller is a function that typically depends on a small set of $(h_{n-k},\varepsilon_{n-k}^{*})$ pairs and the asymptotic order of accuracy for the method $p$, i.e.,
$\tilde{h}=\mathcal{C}(h_{n},\varepsilon_{n}^{*},h_{n-1},\varepsilon_{n-1}^{*},\ldots,p).$
There are a myriad of choices for controllers $\mathcal{C}$ in the literature, but the simplest approach is the so-called I controller. Since $y_{n+1}$ has order $p$, then from [7, Sec. II.3] there exists a function $c(t)$ independent of $h_{n}$ such that,
@@
$\Rightarrow\quad$
@@
Since we have just computed the estimate $\|\varepsilon_{n}^{*}\|$ based on a step with size $h_{n}$, a piecewise constant approximation $c(t)\approx c_{n}$ allows us to “solve” for $\|c_{n}\|=\|\varepsilon_{n}^{*}\|/h_{n}^{p+1}$. Assuming that $c(t)$ will remain relatively constant from one step attempt to the next, the candidate step size $\tilde{h}$ is computed to exactly attain an error goal of 1 (or any toltol):
$tol=\|c_{n}\|\tilde{h}^{p+1}\quad\Rightarrow\quad\tilde{h}=h_{n}\left(\frac{tol}{\|\varepsilon_{n}^{*}\|}\right)^{1/(p+1)}.$
This formula automatically grows or shrinks the step size based on whether the error test $\|\varepsilon_{n}^{*}\|^{1/(p+1)}\leq tol$ passed or failed, respectively. Due to the multiple assumptions above, a safety factor $\sigma<1$ is typically applied,
$\tilde{h}=\sigma h_{n}\left(\frac{tol}{\|\varepsilon_{n}^{*}\|}\right)^{1/(p+1)}.$
More advanced approaches for $\mathcal{C}$ are typically based on control theory, and use additional $(h_{n-k},\varepsilon_{n-k}^{*})$ values to build higher-degree piecewise polynomial approximations of the principal error function [10, 14, 15, 11, 12, 13].
@@ 3 Multirate Temporal Controllers
Multirate temporal adaptivity has not been widely studied in the literature. The two articles we are aware of are [16] and [17]. The former is specific to “MrGARK” methods [18], that require a more rigid algorithmic structure than the MRI algorithm introduced in Section 1. The latter works with MRI methods, and will serve as a baseline comparison for our work. In the following subsections, we consider three such approaches.
@@ 3.1 Coupled Stepsize ($H\text{-}h$”) Control
We note, however, that after introducing these controllers in [17], we found that they struggled on applications where the optimal “multirate ratio” $M_{n}=H_{n}/h_{n}$ was large. Specifically, we found that when $M_{n}\leq 10$ these controllers performed well (consistent with the test problems in [17]), but their performance deteriorated for $M_{n}\geq 20$. In the following subsection, we briefly explain this phenomenon.
3.1.1 Asymptotic analysis of $H\text{-}h$ controller limitations
As an exemplar of the $H\text{-}h$ controller family, we analyze the Constant-Constant H-h Controller (MRI-CC) introduced in [17] in the limit of large multirate ratio. We note that MRI-CC mimics the I controller from Section 2, but updates $M_{n}$ using a formula that couples the fast and slow error estimates from the most recent time step,
$\begin{split}H^{CC}_{n+1}&=H_{n}\|\varepsilon^{\mathfrak{s}}_{n}\|^{-\frac{k_{1}}{P}}\\ M^{CC}_{n+1}&=M_{n}\|\varepsilon^{\mathfrak{s}}_{n}\|^{-\frac{(p+1)k_{1}}{Pp}}\|\varepsilon^{\mathfrak{f}}_{n}\|^{\frac{k_{2}}{p}},\end{split}$ (6)
where $k_{1}$ and $k_{2}$ are controller parameters, $P$ and $p$ are the local orders of accuracy of the embeddings for the slow and fast solvers, and $\varepsilon^{\mathfrak{s}}_{n}$ and $\varepsilon^{\mathfrak{f}}_{n}$ the estimated local errors for the time step $t_{n}\to t_{n}+H_{n}$, respectively.
By comparison, if separate I controllers were performed independently at each time scale, the corresponding multirate ratio would be
$M^{I}_{n+1}=M_{n}\|\varepsilon^{\mathfrak{s}}_{n}\|^{-\frac{1}{P}}\|\varepsilon^{\mathfrak{f}}_{n}\|^{\frac{1}{p}}.$ (7)
To understand the limiting case, it suffices to consider a problem with decoupled fast and slow time scales with $\|\varepsilon^{\mathfrak{s}}_{n}\|=1$ and $\|\varepsilon^{\mathfrak{f}}_{n}\|\gg 1$, which should result in $H_{n+1}=H_{n}$ and $M_{n+1}\gg M_{n}$. Taking the limit as $\|\varepsilon^{\mathfrak{f}}_{n}\|\to\infty$ of the ratio $M_{n+1}^{CC}/M_{n+1}^{I}$, we have
$\lim_{\|\varepsilon_{n}^{\mathfrak{f}}\|\to\infty}\frac{M_{n+1}^{CC}}{M_{n+1}^{I}}$ $=\|\varepsilon_{n}^{\mathfrak{s}}\|^{-\frac{1}{P}\left(\frac{(p+1)k_{1}}{p}-1\right)}\lim_{\|\varepsilon_{n}^{\mathfrak{f}}\|\to\infty}\|\varepsilon_{n}^{\mathfrak{f}}\|^{\frac{k_{2}-1}{p}}.$
The parameters $k_{1}=0.42$ and $k_{2}=0.44$ are determined in [17] by optimizing controller performance across a suite of test problems. Inserting these values and noting that $p>0$, we see that
$\lim_{\|\varepsilon_{n}^{\mathfrak{f}}\|\to\infty}\frac{M_{n+1}^{CC}}{M_{n+1}^{I}}=\|\varepsilon_{n}^{\mathfrak{s}}\|^{-\frac{1}{P}\left(\frac{0.42(p+1)}{p}-1\right)}\lim_{\|\varepsilon_{n}^{\mathfrak{f}}\|\to\infty}\|\varepsilon_{n}^{\mathfrak{f}}\|^{-\frac{0.56}{p}}=0,$
indicating that although the physical problem may necessitate large $M_{n+1}$ values (and thus very small fast steps $h_{n}$), the MRI-CC controller will artificially limit this value, potentially leading to computational inefficiency or inaccuracy. This effect will be readily visible in the numerical results in Section 5 below.
@@ 3.2 Decoupled” Multirate Control
where $P$ and $p$ are the orders of the MRI method and inner solver, respectively. Here, $(H_{k},\varepsilon_{k}^{\mathfrak{s}})$ are the stepsize and local error estimates for time step $k$ at the slow time scale, and $(h_{k,\ell},\varepsilon_{k,\ell}^{\mathfrak{f}})$ are the stepsize and local error estimates for the fast substep $\ell$ within the slow step $k$. We note that in this approach, both $\mathcal{C}^{\mathfrak{s}}$ and $\mathcal{C}^{\mathfrak{f}}$ are decoupled, and thus selection of $\tilde{H}$ and $\tilde{h}$ occurs independently through application of any pair of single-rate controllers. We summarize their use in the following pseudocode.
@@
    Let: $y_{n+1}^{*}=Y_{s}$, $\tilde{y}_{n+1}=\tilde{v}_{s}(\theta_{F,s})$, and $\varepsilon_{n}^{\mathfrak{s}}=\tilde{y}_{n+1}-y^{*}_{n+1}$.
@@ 3.3 Stepsize-tolerance ($H\text{-}Tol$”) Control
Our second proposed family are the so-called $H\text{-}Tol$” multirate controllers. As outlined in Section 2, standard controllers attempt to adapt step sizes to control local error within each step to achieve a requested tolerance. However, MRI methods must ask another “inner” adaptive fast-scale solver to produce the stage solutions $v_{i}(\theta_{F,i})$, that result from numerically integrating over each interval $\theta_{0,i},\theta_{F,i}]$. Local errors within the fast integrator may accumulate, resulting in an overall fast-solver error $\varepsilon^{\mathfrak{f}}_{i,true}$ that could exceed the requested tolerance. If that inner solver can produce both $v_{i}(\theta_{F,i})$ and an estimation of the accumulated error, $\varepsilon^{\mathfrak{f}}_{i}$, then the tolerances provided to that inner solver can be adjusted to ensure MRI stage solutions $Y_{i}=v_{i}(\theta_{F,i})$ that are within the overall tolerances requested of the outer MRI method.
@@
To this end, first assume that at step $n$, we define $\text{tolfac}_{n}$ as a multiplicative factor indicating how much more accurate the inner solver should be than the MRI method, i.e., if the MRI method strives to achieve local errors $\|\varepsilon_{n}^{\mathfrak{s}}\|\leq 1$, then the inner solver strives to achieve local substep errors $\|\varepsilon_{n,m}^{\mathfrak{f}}\|\leq\text{tolfac}_{n}$. Additionally assume that the inner solver is run in “local extrapolation” mode [7, Sec. II.7], i.e., the solution is one order more accurate than the embedding and thus $\|\varepsilon_{n,m,true}^{\mathfrak{f}}\|\approx h_{n,m}\|\varepsilon_{n,m}^{\mathfrak{f}}\|$. Then accounting for accumulation of temporal error in the inner solver, the overall fast error over the slow step $t_{n},t_{n}+H_{n}]$ has the form
$\|\varepsilon^{\mathfrak{f}}_{n}\|\leq\chi_{n}H_{n}\text{tolfac}_{n},$ (9)
where $\chi_{n}$ is independent of $\text{tolfac}_{n}$, but may vary in time. We prove this result in Lemma 3.1, after defining appropriate notation. At stage $i$ of slow step $n$, the inner solver approximates solutions to
$v_{i}^{\prime}(\theta)=f^{\mathfrak{f}}(\theta,v_{i}(\theta))+r_{i}(\theta),\quad v_{i}(\theta_{0,i})=v_{0,i},\quad\theta\in[\theta_{0,i},\theta_{F,i}],$ (10)
where the forcing function $r_{i}(\theta)=r(\theta;t_{i},y_{i})$ is constructed from slow-stage data at preceding stages $j<i$ in the step. Let $v_{i,m}$ denote the numerical approximation produced by the adaptive inner solver at the inner nodes $\theta_{m}$ (with $\theta_{0}=\theta_{0,i}$ and $\theta_{M}=\theta_{F,i}$), i.e., $v_{i,m}\approx v_{i}(\theta_{m})$ for
$\theta_{0,i}=\theta_{0}<\theta_{1}<\cdots<\theta_{m}<\cdots<\theta_{M}=\theta_{F,i},$
with substep sizes $h_{m}=\theta_{m+1}-\theta_{m}$. We assume the inner solver is a one-step method, so that
$v_{i,m+1}=v_{i,m}+h_{m}\Psi^{\mathfrak{f}}_{h_{m}}(\theta_{m},v_{i,m})=:\Phi_{h_{m}}(\theta_{m},v_{i,m}),\quad m=0,\dots,M-1,$ (11)
for some one-step method $\Phi_{h}$ applied to (10). Here, $\Psi^{\mathfrak{f}}_{h_{m}}(\theta_{m},v_{i,m})$ approximates $f^{\mathfrak{f}}(\theta,v_{n}(\theta))+r_{i}(\theta)$ (possibly using multiple internal stages).
Denote the global and local errors after $m$ steps respectively as
$e^{\mathfrak{f}}_{i,m}:=v_{i,m}-v_{i}(\theta_{m}),\quad\varepsilon_{i,m}^{\mathfrak{f}}:=\Phi_{h_{m}}\big(\theta_{m},v_{i}(\theta_{m})\big)-v_{i}(\theta_{m+1}).$ (12)
Lemma 3.1.
Assume that $f^{\mathfrak{f}}(\theta,v)$ is locally Lipschitz continuous in $v$, uniformly in $\theta$ with constant $L_{f}$ and that the inner error controller satisfies the local bound
$\|\varepsilon_{i,m,true}^{\mathfrak{f}}\|\leq h_{m}\text{tolfac}_{n},\qquad m=0,\dots,M-1.$ (13)
Then the accumulated fast error at the end of the inner solve (11) satisfies
$\|e^{\mathfrak{f}}_{i,m}\|\leq\chi_{i}\,(\theta_{F,i}-\theta_{0,i})\,\text{tolfac}_{n},$ (14)
where the constant $\chi_{i}$ is independent of $\text{tolfac}_{n}$.
In particular, if the MRI method satisfies $\sum_{i=1}^{s}\left(\theta_{F,i}-\theta_{0,i}\right)=cH_{n}$ for a fixed constant $c$, then the accumulated error over the $s$ fast IVP solves satisfies
$\|\varepsilon_{n}^{\mathfrak{f}}\|\leq\chi_{n}H_{n}\text{tolfac}_{n},$ (15)
where $\chi_{n}=c\max_{i}\chi_{i}$.
Proof.
First, we observe that
$e^{\mathfrak{f}}_{i,m+1}$ $=v_{i,m+1}-v_{i}(\theta_{m+1})$
$=\Phi_{h_{m}}(\theta_{m},v_{i,m})-\Phi_{h_{m}}\big(\theta_{m},v_{i}(\theta_{m})\big)+\Phi_{h_{m}}\big(\theta_{m},v_{i}(\theta_{m})\big)-v_{i}(\theta_{m+1})$
$=\Phi_{h_{m}}(\theta_{m},v_{i,m})-\Phi_{h_{m}}(\theta_{m},v_{i}(\theta_{m}))+\varepsilon_{i,m}^{\mathfrak{f}}.$
Under the assumption that $f^{\mathfrak{f}}(\theta,v)$ has the Lipschitz constant $L_{f}$, it follows from (11) that $\Phi_{h}(\theta,v)$ also satisfies the Lipschitz condition (in its second argument) with the Lipschitz constant $L_{\Phi_{h_{m}}}=1+h_{m}L_{f}$. Therefore,
$\|e^{\mathfrak{f}}_{i,m+1}\|\leq(1+L_{f}h_{m})\|e^{\mathfrak{f}}_{i,m}\|+\|\varepsilon_{i,m}^{\mathfrak{f}}\|.$
Solving this recursion, applying a discrete Grönwall inequality, and using the fact that $\sum_{m=0}^{M-1}h_{m}=\theta_{F,i}-\theta_{0,i}$, yields
$\|e^{\mathfrak{f}}_{i,M}\|\leq\exp\Bigl(L_{f}\sum_{m=0}^{M-1}h_{m}\Bigr)\sum_{m=0}^{M-1}\|\varepsilon_{i,m}^{\mathfrak{f}}\|=\exp\bigl(L_{f}(\theta_{F,i}-\theta_{0,i})\bigr))\sum_{m=0}^{M-1}\|\varepsilon_{i,m}^{\mathfrak{f}}\|.$
Using (13) gives
$\|e^{\mathfrak{f}}_{i,M}\|\leq\exp\bigl(L_{f}(\theta_{F,i}-\theta_{0,i})\bigr)\sum_{m=0}^{M-1}h_{m}\text{tolfac}_{n}\ =\chi_{i}\,(\theta_{F,i}-\theta_{0,i})\,\text{tolfac}_{n}$
with $\chi_{i}=\exp\bigl(L_{f}(\theta_{F,i}-\theta_{0,i})\bigr)$. This proves (14), and consequently, (15). ∎
@@
To convert equation (9) into a multirate controller, we replace the inequality with an equality, and then regroup terms as $\|\varepsilon^{\mathfrak{f}}_{n}\|=\left(\chi_{n}H_{n}\right)\left(\text{tolfac}_{n}\right)^{0+1}$, we see that this fits the standard asymptotic error assumption used in single-rate controllers (5), through identifying the control parameter $\text{tolfac}_{n}$ with $h_{n}$, $\chi_{n}H_{n}$ with $\|c(t_{n})\|$ (note that this is independent of $\text{tolfac}_{n}$), and by defining the “order” $p=0$. Thus, any single-rate controller that relies on the asymptotic error assumption (9) may be used to adjust $\text{tolfac}_{n}$ between slow step attempts.
@@
    Let: $y^{*}_{n+1}=Y_{s}$, $\tilde{y}_{n+1}=\tilde{v}_{s}(\theta_{F,s})$, $\varepsilon_{n}^{\mathfrak{s}}=\tilde{y}_{n+1}-y^{*}_{n+1}$, and retrieve the accumulated fast error $\varepsilon^{\mathfrak{f}}_{n}$ from the inner solver.